OpenStudy (konradzuse):

In 1965, about 44% of the U.S. adult population had never smoked cigarettes. A national health survey of 1205 U.S. adults (presumably selected randomly) during 2006 revealed that 615 had never smoked cigarettes. Reference: Ref 18-5 A 96% confidence interval for the proportion of U.S. adults in 2006 that have never smoked is A. 0.461 to 0.560. B. 0.470 to 0.551. C. 0.481 to 0.540. D. 0.487 to 0.534.

OpenStudy (konradzuse):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

this sounds familiar...

OpenStudy (konradzuse):

Hmm lets see....

jimthompson5910 (jim_thompson5910):

we did this one?

OpenStudy (konradzuse):

In 1965, about 44% of the U.S. adult population had never smoked cigarettes. A national health survey of 1205 U.S. adults (presumably selected randomly) during 2006 revealed that 615 had never smoked cigarettes. Reference: Ref 18-5 Suppose you wished to test whether there has been a change since 1965 in the proportion of U.S. adults that have never smoked cigarettes. Which of the following are the appropriate hypotheses? A. H0: p = 0.44, Ha: p > 0.44 B. H0: p = 0.51, Ha: p 0.51 C. H0: p = 0.44, Ha: p 0.44 D. H0: = 0.44, Ha: 0.44 Correct Points Earned: 1/1 Your Response: C 4. In 1965, about 44% of the U.S. adult population had never smoked cigarettes. A national health survey of 1205 U.S. adults (presumably selected randomly) during 2006 revealed that 615 had never smoked cigarettes. Reference: Ref 18-5 The P-value of the test of hypotheses in question 3 is A. greater than 0.10. B. between 0.05 and 0.10. C. between .01 and 0.05. D. below 0.01.

OpenStudy (konradzuse):

We did these :0

jimthompson5910 (jim_thompson5910):

oh gotcha

OpenStudy (konradzuse):

good memory :D.

OpenStudy (konradzuse):

I'm going with B!

OpenStudy (anonymous):

B

OpenStudy (anonymous):

b

jimthompson5910 (jim_thompson5910):

nope, not B

OpenStudy (anonymous):

kay dont hate me

OpenStudy (anonymous):

dang it! i thought i had it

OpenStudy (anonymous):

sowwwyy

OpenStudy (konradzuse):

Hi team, thanks for the backup!

OpenStudy (anonymous):

oh for sure!(;

OpenStudy (anonymous):

you got it kboyyy

OpenStudy (konradzuse):

See you later in chat QT pies.

OpenStudy (konradzuse):

Ok back to this stupid stats.... We gotta do some P<Z stuff right?

OpenStudy (konradzuse):

or is this the confidence level BS...?

OpenStudy (anonymous):

im never going back on chat EVER

OpenStudy (anonymous):

me either!!!

OpenStudy (konradzuse):

NormalCfl(Wolfram).

OpenStudy (anonymous):

we got kicked off... hahahahha

OpenStudy (anonymous):

sooooo BA

OpenStudy (konradzuse):

You're cray... What did you say?

OpenStudy (anonymous):

i told someone to die

jimthompson5910 (jim_thompson5910):

sample proportion: m = x/n = 615/1205 = 0.510373 Margin of error: E = C*sqrt((m*(1-m))/n) E = 2.0537489*sqrt((0.510373*(1-0.510373))/1205) E = 2.0537489*sqrt((0.510373*(1-0.510373))/1205) E = 0.02957538 ------------------------------------- 96% Confidence interval: (m - E, m + E) (0.510373 - 0.02957538, 0.510373 + 0.02957538) (0.48079762, 0.53994838) (0.481, 0.540)

OpenStudy (konradzuse):

So It's c....

OpenStudy (konradzuse):

I'm going to write this formula down so I remember it...

OpenStudy (konradzuse):

Okay team onto the next Q!

OpenStudy (anonymous):

okay!